OpenStudy (anonymous):

Write the equation of the line containing the altitude ... Write the equation of the line containing the median that passes through point C in slope-intercept form .. Points are a (-4,-2) b (4,4) C 18,-8).

OpenStudy (jdoe0001):

@Earlgirl have you covered slopes yet?

OpenStudy (anonymous):

Yes

OpenStudy (jdoe0001):

ok have you plotted those points yet?

OpenStudy (jdoe0001):

do you know where they go?

OpenStudy (anonymous):

I had three questions regarding these these points. The first question asked me write an equation of the line containing th altitude that passes through B in standard form. I first did x1+x2/2 and y1+y2/2 to get mx1 and mx2. Then I found the slope (m) . I plugged that into y-y1=m(x-x1)

OpenStudy (anonymous):

The second question asked me to write an equation for the line containing perpendicular bisector of AC in point slope form . I followed th same process as I did with the median

OpenStudy (jdoe0001):

right , notice the graph

OpenStudy (jdoe0001):

that's correct, the blue line, coming down from point B, is the indeed the altitude because is the distance for the widest GAP in the triangle, thus the altitude the purple line, would be the median through point "c" because is coming from the MIDDLE of the opposite side of "c" and passes through "c"

OpenStudy (anonymous):

So would I repeat the as,e process to find the equation for the median ?

OpenStudy (jdoe0001):

well, you see, dunno what you've done so far... to find the altitude, what you're really being asked is "what is the equation of a line that is PERPENDICULAR to line AC and passes through point B" so you'd need to firstly find the equation of the line AC then keeping in mind that a PERPENDICULAR line to that equation will have a NEGATIVE RECIPROCAL slope use that slope and the point B (4,4) to get its equation

OpenStudy (jdoe0001):

notice the picture, if we straighten up the triangle and make AC the base, then the altitude will be PERPENDICULAR to AC and will pass through point B

OpenStudy (anonymous):

Y+5=-6/22x-7 or y+5=22/6x-7

OpenStudy (anonymous):

For the altitude y=3/4x+1

OpenStudy (jdoe0001):

now the 2nd part is quite simple you're asked for a line that comes out from the median and goes through "c" so for that, you'd just need to get the MidPoint formula to get the MidPoint of AB and then once you have that point, you'd want the equation of a line that goes through THAT POINT and "c"

OpenStudy (anonymous):

X1+x2/2 and y1+y2/2

OpenStudy (jdoe0001):

for the midpoint, yes

OpenStudy (anonymous):

What do I use for the other set of points ?

OpenStudy (anonymous):

C = 18,-8

OpenStudy (jdoe0001):

y=3/4x+1 <--- I didn't get that btw

OpenStudy (jdoe0001):

C = 18,-8 <-- yes

OpenStudy (jdoe0001):

hmmm lemme recheck myself

OpenStudy (jdoe0001):

yea... I didn't get y=3/4x+1

OpenStudy (anonymous):

Hmmm what did you get? I used points (-4,-2)(4,4)

OpenStudy (anonymous):

I got zero and 1 as the midpoints

OpenStudy (jdoe0001):

(0, 1) is correct for the midpoint of AB

OpenStudy (jdoe0001):

so the purple line will be the "equation of a line that passes through (0,1) and (18, -8)"

OpenStudy (anonymous):

Ok thanks . So basically you follow the same process for find the equation for each line

OpenStudy (jdoe0001):

yes, using 2 points, getting the slope and and then plugging the values in the "point-slope form"

OpenStudy (jdoe0001):

http://mathb.in/12767 <--- for the first one

OpenStudy (anonymous):

I forgot to simplify